OpenStudy (simplixity):

A water balloon is launched at a speed of 24 m/s and an angle of 36 degrees above the horizontal. The water balloon hits a tall building located 26 m from the launch pad. At what height above the ground level will the water balloon hit the building? Calculate the answer in meters (m) and rounded to three significant figures

jhonyy9 (jhonyy9):

do you can drawing it please this case ? i think will be usefully

OpenStudy (simplixity):

so far I have this drawn but I'm not sure if it's correct.. 

jhonyy9 (jhonyy9):

so this 26 m i think not will be this imagined by you there what mean 26 m from the launch pad ?

OpenStudy (mhchen):

Lol, kinematic equations and 2d problems. Physics I bet. So your launch angle is 36 degrees, and velocity is 24 m/s. That is correct. Forget the 24cm part for now. What you need to do first is to divide the 24m/s into x and y components. Do you know decomposition of vectors?

OpenStudy (mhchen):

Basically dividing the 24m/s into x m/s and y m/s

OpenStudy (mhchen):

And then after that, you have to find the time the water balloon reaches 24m in the x-direction using the x-direction velocity component.

OpenStudy (simplixity):

which equation would I use for that?

OpenStudy (mhchen):

So to divide the 24m/s into components. X-direction: cos(36) * 24 That gives you it's speed in the x-direction only. Get it?

OpenStudy (simplixity):

oh! so then sin(36)*24 for y direction, right?

OpenStudy (mhchen):

Yes. We'll worry about that later though.

OpenStudy (simplixity):

I got 19.42 for the speed in x-direction

OpenStudy (mhchen):

Alrighty. So then once you have the x-velocity, and assuming there's no air-friction. We use \[x = x_{0} + v_{0}t + \frac{ 1 }{ 2 }at ^{2}\] You were taught this formula right?

OpenStudy (simplixity):

yeah, I recognize it. is it okay if I work on it and then you check my answers?

OpenStudy (mhchen):

suuure

OpenStudy (simplixity):

I got t=3.96

OpenStudy (simplixity):

@mhchen

OpenStudy (mhchen):

Erm hold on. so x = 0 + vt + 0 24 = (19.42)t t = (24/19.42) t = 1.24s I got something different :\

OpenStudy (mhchen):

a = 0 and initial position = 0

OpenStudy (simplixity):

oh, there's no gravity?

OpenStudy (mhchen):

that's only for y-direction.

OpenStudy (simplixity):

oh okay

OpenStudy (mhchen):

Okay so now that you've got t. You can do the same equation for y: \[y = y_{0} + v_{0}t + \frac{ 1 }{ 2 }at ^{2}\] And this time you would put the gravity in.

OpenStudy (simplixity):

and I'm using what I got for sin right? or am I still using the cos

OpenStudy (mhchen):

you use sin this time. sin finds the y-component.

OpenStudy (simplixity):

I got 9.95

OpenStudy (mhchen):

yeah same here.

OpenStudy (mhchen):

I'm in physics as well. I remember doing these problems like last week. It gets more complicated later on :(

OpenStudy (simplixity):

oh gosh.. I hate it. I'm confused like half the time :( so would this be the height then?

OpenStudy (mhchen):

Yeah.

OpenStudy (simplixity):

thanks for your help, I appreciate it!

OpenStudy (mhchen):

np